# LLM-Based Intent Processing and Network Optimization Using Attention-Based Hierarchical Reinforcement Learning

### v1

@@ Abstract
Intent-based network automation is a promising tool to enable easier network management however certain challenges need to be effectively addressed. These are: 1) processing intents, i.e., identification of logic and necessary parameters to fulfill an intent, 2) validating an intent to align it with current network status, and 3) satisfying intents via network optimizing functions like xApps and rApps in O-RAN. This paper addresses these points via a three-fold strategy to introduce intent-based automation for O-RAN. First, intents are processed via a lightweight Large Language Model (LLM). Secondly, once an intent is processed, it is validated against future incoming traffic volume profiles (high or low). Finally, a series of network optimization applications (rApps and xApps) have been developed. With their machine learning-based functionalities, they can improve certain key performance indicators such as throughput, delay, and energy efficiency. In this final stage, using an attention-based hierarchical reinforcement learning algorithm, these applications are optimally initiated to satisfy the intent of an operator. Our simulations show that the proposed method can achieve at least $12\%$ increase in throughput, $17.1\%$ increase in energy efficiency, and $26.5\%$ decrease in network delay compared to the baseline algorithms.
@@
Intent-based network automation, Attention-based hierarchical reinforcement learning, network optimization
@@ I Introduction
Achieving intent-based automation can be highly challenging in modern-day 5G paradigms like Open Radio Access Network (O-RAN). Instead of tweaking specific parameters through vendor-specified interfaces, Mobile Network Operators (MNOs) can now express their intentions in natural language. This approach introduces a huge challenge in the telecommunication industry since intents can vary widely based on network conditions, customer demands, and industry-specific terminology [1]. Therefore, a contextual understanding of the intents is vital for accurate intent processing. Large Language Models (LLMs) excel at capturing such context and understanding the underlying meaning of sentences (intents) within a broader spectrum. For instance, Wang et al. propose a Transformer model based on an LLM where various forms of network data (packet size, time stamps, and end-to-end delay) are integrated into a unified feature space for refining network control policies to align with an MNO’s intent [2]. However, modern-day communication systems require continuous monitoring of the Key Performance Indicators (KPIs) to serve users having different kinds of Quality-of-Service (QoS) requirements. Existing LLM-based approaches like [2] are not specially tailored to extract crucial information associated with what performance metric the operator is aiming to improve and by how much.
@@
Processing intents expressed in natural language by the MNO is crucial for deriving policy for network performance optimization. Accurate processing of intents can provide MNOs with feedback on how their intents might impact crucial performance metrics like energy efficiency or throughput. For example, an intent: “Increase energy efficiency by $30\%$” during peak hours can degrade performance due to high traffic volume. This kind of scenario requires a sophisticated method of intent validation using contemporary network status to ensure a seamless user experience without compromising any QoS.
The introduction of paradigm-shifting concepts like O-RAN paves the way for deploying network-optimizing applications known as xApps and rApps in RAN Intelligent Controllers (RICs). These apps can be initiated based on MNO’s processed intents. Based on this idea, in [1], intents are converted into goals for a Hierarchical Reinforcement Learning (HRL) algorithm to initiate xApps or rApps for intent fulfillment. This decision-making process is computationally intensive, given the vast number of potential rApp or xApp combinations that can optimize performance. An AI agent, therefore, must not only choose from numerous options but also identify the most appropriate combination of network-optimizing apps that align with an MNO’s intent.
@@
To this end, we propose a three-fold strategy to perform intent processing, validation, and RIC application initiation based on intents. First, we process intents using Bidirectional Encoder Representations from Transformers (BERT) which is an LLM model that can be fine-tuned for specific tasks like intent extraction, which allows it to learn task-specific patterns of the data [3]. In the next step, intent validation is performed using a Transformer-based time series predictor for upcoming traffic volume [4]. This ensures that the intended adjustments to the network are not only theoretically sound but also practical and feasible given expected future traffic patterns. Lastly, to overcome the computational burden of selecting the appropriate combination of xApps and rApps from a high number of possible choices, we utilize an attention-based HRL framework that takes the processed intents (magnitude of the intended performance metric by an MNO) as goals. By filtering out only the feasible options [5], via an attention mechanism, the system can significantly reduce the computational burden
@@
Unlike the previous works, the proposed scheme provides complete intent-based automation from an O-RAN perspective via intent processing, validation, and HRL-based performance optimization using xApps and rApps. The simulation results demonstrate that the proposed scheme yields significant improvements compared to two baseline approaches: HRL without attention and intent validation, and a single-application scenario developed using Deep RL (DRL). The proposed approach surpasses the HRL baseline in terms of throughput, delay, and energy efficiency by $12.02\%$, $26.5\%$, and $17.1\%$ respectively. Moreover, compared to the DRL baseline, the proposed method achieves superior results across the same performance metrics, with improvements of $17.25\%$, $48.6\%$, and $39.3\%$ respectively.
The organization of this manuscript is outlined as follows: Section II offers a short review of the existing literature associated with this research. This is succeeded by Section III, which elaborates on the system model in depth. Section IV is specifically devoted to elaborately presenting our proposed approach. Performance assessment and a comparative examination of our approach against the baseline algorithms are detailed in Section V. Lastly, Section VI presents the concluding remarks of the paper.
@@ II Related work
Even though using LLMs for intent processing for network optimization is a relatively new concept, a few works have been found in the literature. Wang et al.’s development of a transformer model based on an LLM is one such example, where multi-modal representation learning is employed to integrate various forms of network data into a unified feature space [2]. Similarly, Kristina et al. have introduced a pipeline that utilizes an LLM to process intents into structured, policy-based abstractions, linking them to APIs for execution [6].
To manage RIC applications based on operator intents, an approach based on HRL is proposed in [7] where intents are processed into goals for the control algorithm to optimize performance. Polese et al. propose a framework to collect control requests (intents) from an MNO and select the optimal Machine Learning (ML) models to achieve the operator’s goals to avoid conflicts [8].
Compared with existing literature, the main contribution of this work lies in deploying an LLM to interpret MNO’s intent and extract specific parameters for further network performance optimization. Processed intents are validated against future wireless traffic conditions. Intents lead HRL agents to initiate and orchestrate multiple network functions with an efficient exploration of action space via an attention mechanism.
@@ III System Model
A cellular system based on O-RAN architecture employing downlink orthogonal frequency division multiplexing is considered in this work. There are $B$ base stations that concurrently serve $U$ users. Within the macro cell’s coverage area, multiple small cells are deployed. The system supports $K$ classes of traffic, and users are connected through dual connectivity across multiple Radio Access Technologies (RATs). These RAT classes, denoted as $R_{1},R_{2},\ldots,R_{n}$, can facilitate various technologies like LTE, 5G.
@@
The system model, depicted in Figure 1, incorporates RAN Intelligent Controllers (RICs), comprising Non-Real-Time (non-RT)-RIC and Near-Real-Time (Near-RT)-RIC. These controllers host both rApps and xApps, focusing on control and optimization tasks across different time scales. We deploy BSs capable of switching bands from 3.5 GHz to mmWave frequencies. This deployment enables us to cater to high-throughput traffic by employing intelligent beamforming techniques. An xApp can be developed to regulate power based on UE location, utilizing minimal transmission power for enhanced energy efficiency. The system facilitates analog beamforming, with each BS employing a uniform linear array of $\mu$ antennas [9]. Beamforming weights for each beamforming vector are implemented using constant modulus phase shifters. Additionally, we assume the presence of a beam steering-based codebook $F$, from which beamforming vectors are selected [9]. Each BS $b$ operates with a transmit power $P_{TX,b}\in P$, where $P$ represents the set of candidate transmit powers. The energy consumption model for the BS is obtained from [10].
Figure 1: System model designed with macro cell and small cells based on O-RAN architecture.
III-A Problem Formulation
@@
Based on the system and network model discussed so far, we introduce multiple xApps and rApps. There are two sets of these RIC applications. The first set consists of intent translation, validation, and meta-controller rApps and the second set consists of traffic steering, beamforming, cell sleeping, power allocation, and handover management applications that can be directly controlled by the first set.
@@
Let us define $\Psi$ as the collection of xApps, rApps, and their various combinations. Let $\sigma$ represent the subset of $\Psi$ which contains at least one element (either an xApp or rApp) that is capable of enhancing network performance in response to operator intents. We also define $\rho$ as the set of potential KPIs that an application might improve, and $Q_{s}$ as the set of QoS requirements that must be met by the system. Given these definitions, the challenge of orchestrating RIC applications based on operator intent can be formulated as follows:
@@
$\begin{split}\max\sum_{\rho\in\varrho}\sum_{q_{s}\in Q_{s}}(C_{\rho}-\varpi\gamma_{q_{s}}),\quad\quad\quad\\ \text{s.t.}\quad\forall(\Psi)\exists(\sigma):\zeta(O)=1,\quad\quad\quad\end{split}$ (1)
@@
where $C$ is the magnitude of the intended performance metric the operator intends to improve. $\varpi$ is the penalty parameter for QoS requirement violation, and $\gamma_{q_{s}}$ represents the number of UEs whose QoS requirements have been violated. The evaluation is binary: a UE is counted only if its QoS standards are not met. QoS violations are assessed based on the QoS parameters associated with each UE’s traffic class. Lastly, $\zeta(O)$ is the proposition that “An xApp can improve a performance metric”, which is either ‘0’ or ‘1’.
IV Proposed Intent Processing and Network Optimization Scheme
In this section, first, we discuss processing intents using an LLM followed by intent validation. The last part of this section consists of an HRL-based attention mechanism for RIC application initiation and orchestration.
IV-A Intent Processing using LLMs
Intent-based networking simplifies network and administrative tasks by automating system operations through the comprehension and execution of operator-defined intents, such as “Increase overall energy efficiency by 10%”, “Boost system throughput by 15%”, or “Reduce network delay by 13%”. These intents focus on improving metrics like throughput, energy efficiency, and network delay. We developed a dataset containing various intents related to 5G and O-RAN performance metrics and annotated it with details like action direction (increase or decrease), change percentage, and targeted metric.
We utilized a lightweight version of BERT, known as ALBERT, for efficient intent processing. This model is particularly suitable for real-time inference in O-RAN scenarios due to its balance of performance and computational efficiency [11]. Using pre-trained ALBERT, we trained our system to predict the specified elements from intent expressions. This intent processing functionality is presented as User Interface (UI) and intent translation rApp in Fig. 2. The job of this rApp is to work as a UI for an MNO and process the intents with crucial information to feed the intent validation rApp and the meta-controller rApp.
IV-B Intent Validation Using Transformer
@@
The second step of this work consists of intent validation. QoS parameters deviate from the originally defined QoS metrics over time due to performance degradation in the wireless networks causing QoS drifts. By deviation, we mean when a QoS parameter is having lesser value than the required. For a specific traffic class, $k$ having a set of QoS requirements $Q_{s}$ and associated performance metrics $C$, we can formulate QoS drift as follows:
$\begin{split}\min\sum_{q_{j}\in Q_{s}}\sum_{c\in C}\{D_{QoS}-A_{QoS}\},\quad\quad\\ \text{s.t.}\quad A_{QoS}\leq D_{QoS}\quad\text{and}\quad A_{QoS}>0.\end{split}$ (2)
@@
In eq.2, $A_{QoS}$ represents the achieved value of a metric (e.g. delay, throughput) for a certain traffic type at time $t$ and $D_{QoS}$ represents the pre-defined value of the same parameter which has to be maintained for optimal performance. The objective here is to validate intents with the assurance that they will not result in QoS drifts. The condition, $A_{QoS}\leq D_{QoS}$ represents the fact that this objective function presented in eq. 2 is only applicable for scenarios when a QoS parameter has a downfall over time compared to the pre-defined value. For cases where $A_{QoS}>D_{QoS}$, we consider there is no QoS drift. If there is no QoS drift for an intent, we consider them automatically valid. If not we calculate the QoS drift, $I_{Q}=D_{QoS}-A_{QoS}$ based on eq. 2.

### v2

@@ Abstract
Intent-based network automation is a promising tool that enables easier network management; however, certain challenges must be addressed effectively. These are: 1) processing intents, i.e., identification of logic and necessary parameters to fulfill an intent, 2) validating an intent to align it with current network status, and 3) satisfying intents via network optimizing applications. This paper addresses these points via a three-fold strategy to introduce intent-based automation for modern 5G architectures. First, intents are processed via a lightweight Large Language Model (LLM). Secondly, once an intent is processed, it is validated against future incoming traffic volume profiles (high or low). Finally, a series of network optimization applications has been developed. With their machine learning-based functionalities, they can improve certain key performance indicators such as throughput, delay, and energy efficiency. In the final stage, using an attention-based hierarchical reinforcement learning algorithm, these applications are optimally initiated to satisfy the intent of an operator. Our simulations show that the proposed method can achieve at least a $12\%$ increase in throughput, a $17.1\%$ increase in energy efficiency, and a $26.5\%$ decrease in network delay compared to the baseline algorithms.
@@
Attention-based hierarchical reinforcement learning, intent-based network automation, network optimization
@@ I Introduction
An intent can be defined as the expectations, including requirements, goals, and constraints for a specific service or network management workflow[1]. It is typically understandable by humans and must be interpreted by machines without ambiguity. Furthermore, it emphasizes “What” needs to be achieved, rather than focusing on “How” to achieve the outcomes, specifying the desired metrics without detailing the process to attain them [1]. It is anticipated that Mobile Network Operators (MNOs) will leverage advancements in natural language processing to express intents in plain language for automated network management. In modern 5G networks, achieving such intent-based automation can be challenging especially because intents can vary widely based on network conditions, customer demands, and industry-specific terminologies [1]. Therefore, a contextual understanding of the intents is vital for accurate intent processing. Large Language Models (LLMs) with its transformer architecture excel at capturing such context and understanding the underlying meaning of sentences (intents) within a broader spectrum [2].
@@
Processing intents expressed in natural language by the human network operator is important to derive policies for network performance optimization. However, allowing any kind of intent to put an impact on the network configurations can lead to serious degradation in performance. For example, an intent: “Increase energy efficiency by $30\%$” during peak hours can degrade performance due to high incoming traffic volume. This kind of scenario requires a sophisticated method of intent validation using contemporary network status to ensure a seamless user experience without compromising any QoS.
The concept of an Open Radio Access Network (Open RAN) facilitates the idea of disaggregating the traditional RAN and promoting interoperability between components from different vendors. This concept enables the deployment of network-optimizing applications (apps) in intelligent RAN controllers, which can be triggered by processed intents. In [3], intents are transformed into goals for a Hierarchical Reinforcement Learning (HRL) algorithm to initiate such applications for intent fulfillment. However, the computational intensity of selecting optimal combinations from numerous applications becomes significant for an AI agent.
@@
To this end, we propose a three-fold strategy to perform intent processing, validation, and application initiation based on intents. We begin by processing intents using a pre-trained LLM named Bidirectional Encoder Representations from Transformers (BERT) [4]. In the next step, intent validation is performed using a Transformer-based time series predictor for upcoming traffic volume [5]. This ensures that the intended adjustments to the network are not only theoretically sound but also practical and feasible given expected future traffic patterns. Lastly, to overcome the computational burden of selecting the appropriate combination of applications from a high number of possible choices, we utilize an attention-based HRL framework that takes the processed intents (magnitude of the intended performance metric by an operator) as goals. By using an attention mechanism to filter only the feasible options [6], the system can significantly reduce the computational burden.
@@
The simulation results demonstrate that the proposed scheme yields significant improvements compared to two baseline approaches: HRL without attention and intent validation, and a single-application scenario developed using Deep RL (DRL). The proposed approach surpasses the HRL baseline in terms of throughput, delay, and energy efficiency by $12.02\%$, $26.5\%$, and $17.1\%$ respectively. Moreover, compared to the DRL baseline, the proposed method achieves superior results across the same performance metrics, with improvements of $17.2\%$, $48.6\%$, and $39.3\%$ respectively.
@@ II Related work
Even though using LLMs to process intents for network optimization is a relatively new concept, a few works have been found in the literature. Wang et al.’s development of a transformer model based on an LLM is one such example, where multi-modal representation learning is employed to integrate various forms of network data into a unified feature space [7]. Similarly, Kristina et al. have introduced a pipeline that utilizes an LLM to process intents into structured, policy-based abstractions, linking them to APIs for execution [2].
To manage network applications based on operator intents, an approach based on HRL is proposed in [3] where intents are processed into goals for the control algorithm to optimize performance. Polese et al. propose a framework to collect control requests (intents) and select the optimal Machine Learning (ML) models to achieve the operator’s goals to avoid conflicts [8].
Compared with existing literature, the main contribution of this work lies in deploying an LLM to interpret a human network operator’s intent and extract specific parameters for further network performance optimization. Processed intents are validated against future wireless traffic conditions. Intents lead HRL agents to initiate and orchestrate multiple network applications with an efficient exploration of action space via an attention mechanism.
@@ III System Model
A cellular system employing downlink Orthogonal Frequency Division Multiplexing (OFDM) is considered in this work. There are $B$ base stations that concurrently serve $U$ users. Within the macro cell’s coverage area, multiple small cells are deployed. The system supports $K$ classes of traffic, and users are connected through dual connectivity across multiple Radio Access Technologies (RATs). These RAT classes can facilitate various technologies like LTE, and 5G.
The system model, shown in Fig. 1, proposes two distinct controllers responsible for managing RAN functionalities: the strategic controller and the tactical controller. The strategic controller on top is responsible for long-term goals, such as network-wide performance optimizations, and policy formulation. It can define the network’s broader objectives, including service quality guarantees, energy efficiency, or overall traffic management for a timespan. The tactical controller on the bottom focuses on real-time decisions, such as handling immediate network changes, performing handovers, or dynamically adjusting resources based on the current network state. It operates within the framework set by the strategic controller but is more focused on near-real-time adjustments to ensure smooth network operation from moment to moment.
@@
The controllers host applications focusing on control and optimization tasks across two different time scales. We deploy BSs capable of switching bands from 5G mid band to high band frequencies. This deployment enables us to cater to high-throughput traffic by employing intelligent beamforming techniques. An application can be developed to regulate power based on UE location, utilizing minimal transmission power for enhanced energy efficiency. The system facilitates analog beamforming, with each BS employing a uniform linear array of $\mu$ antennas [9]. Beamforming weights for each beamforming vector are implemented using constant modulus phase shifters. Additionally, we assume the presence of a beam steering-based codebook $F$, from which beamforming vectors are selected [9]. Each BS $b$ operates with a transmit power $P_{TX,b}\in P$, where $P$ represents the set of candidate transmit powers. The energy consumption model for the BS is obtained from [10].
III-A Problem Definition
@@
Based on the system and network model discussed so far, we introduce multiple applications. There are two sets of them. The first set consists of intent translation, validation, and meta-controller (associated with our RL-based control algorithm). The second set consists of traffic steering, beamforming, cell sleeping, power allocation, and handover management applications that can be directly controlled by the first set. All applications from the first set, along with cell sleeping, operate on a time frame longer than 1 second within the strategic controller, while the others preferably operate on a time scale between 10 ms and 1s within the tactical controller.
@@
Let us define $\Psi$ as the collection of applications, and their various combinations. Let $\sigma$ represent the subset of $\Psi$ which contains at least one element that is capable of enhancing network performance in response to operator intents. We also define $\rho$ as the set of potential KPIs that an application might improve, and $Q_{s}$ as the set of QoS requirements that must be met by the system. Given these definitions, the challenge of orchestrating applications based on operator intent can be formulated as follows:
@@
$\begin{split}P_{1}:\max\sum_{\rho\in\varrho}\sum_{q_{s}\in Q_{s}}(C_{\rho}-\varpi\gamma_{q_{s}}),\quad\quad\quad\\ \text{s.t.}\quad\forall(\Psi)\exists(\sigma):\zeta(O)=1,\quad\quad\quad\end{split}$ (1)
@@
where $C$ is the magnitude of the intended performance metric the operator intends to improve. $\varpi$ is the penalty parameter for QoS requirement violation, and $\gamma_{q_{s}}$ represents the number of UEs whose QoS requirements have been violated. The evaluation is binary: a UE is counted only if its QoS standards are not met. Lastly, $\zeta(O)$ is the proposition that “An application can improve a performance metric”, which is either ‘0’ or ‘1’.
@@
The second problem we want to address in this work is associated with intent validation. QoS parameters deviate from the originally defined QoS metrics over time due to performance degradation in the wireless networks, causing QoS drifts. For a specific traffic class $k$ ($K=\{k_{1},k_{2},\ldots,k_{\eta}\}$), with its set of QoS requirements $Q_{s}$ ($Q_{s}=\{Q_{s_{1}},Q_{s_{2}},\ldots,Q_{s_{j}}\}$) and corresponding performance metrics $C$, the QoS drift can be formulated as follows:
$\begin{split}P_{2}:\min\sum_{q_{s}\in Q_{s}}\sum_{c\in C}\{D_{QoS}-A_{QoS}\},\quad\quad\\ \text{s.t.}\quad A_{QoS}\leq D_{QoS}\quad\text{and}\quad A_{QoS}>0.\end{split}$ (2)
@@
In (2), $A_{QoS}$ represents the achieved value of a metric (e.g. delay, throughput) for a certain traffic type at time $t$, and $D_{QoS}$ represents the pre-defined value of the same parameter which has to be maintained for optimal performance. The objective here is to validate intents that will not result in QoS drifts. For cases where $A_{QoS}>D_{QoS}$, we consider there is no QoS drift and the intent is valid. Otherwise, we calculate the QoS drift, $I_{Q}=D_{QoS}-A_{QoS}$ based on (2).
Another objective of this work is to reduce the large action space, which consists of numerous potential combinations of applications. To solve this problem, we define an attention function $\alpha(s_{t},i,a)$ that scores each action $a\in A$ based on the current state $s_{t}$ and the operator’s intent $i$.
IV Proposed Methodology
In this section, first, we discuss processing intents using an LLM and then predictive intent validation. The last part of this section consists of an HRL-based attention mechanism for application initiation and orchestration.
IV-A Intent Processing using LLM
Intent-based networking simplifies network and administrative tasks by automating system operations through the comprehension and execution of operator-defined intents, such as “Increase overall energy efficiency by 10%”, “Boost system throughput by 15%”, or “Reduce network delay by 13%”. These intents focus on improving metrics like throughput, energy efficiency, and network delay. Our approach uses few-shot learning to classify intents and extract relevant keywords. The methodology involves constructing a prompt that includes a few examples of intents with their corresponding types and keywords. This prompt is then used to query the LLM, which classifies new intents and extracts the required information based on the provided examples. Before processing new intents, a set of examples is prepared. Each example consists of:
  • Intent: A natural language statement describing the intent.
  • Type: The category or classification of the intent.
  • Keywords: Key phrases or terms extracted from the intent.
A prompt is constructed by sequentially adding the prepared examples. Each example is formatted to include the intent ($I=\{i_{1},i_{2},\ldots,i_{n}\}$), type ($I_{t_{y}}=\{t_{y_{1}},t_{y_{2}},...,t_{y_{m}}\}$), and keywords ($\Lambda=\{\lambda_{1},\lambda_{2},\ldots,\lambda_{w}\}$). After adding all the examples, the new intent is appended to the prompt ($p$), followed by a request for classification and keyword extraction. The constructed prompt is sent to the pre-trained LLM (a lightweight version of BERT) [11]. The LLM uses the context provided by the examples to understand the task and generate a response that includes the classification and keywords for the new intent. The response from the LLM is parsed to extract the classified type of the new intent and the extracted keywords. Each new intent’s classified type and keywords are compiled into a result set.
Algorithm 1 Intent Processing Using LLM
1:$E=\{(i_{1},t_{y_{1}},\lambda_{1}),(i_{2},t_{y_{2}},\lambda_{2}),...\}$, $\Xi=\{(i,t_{y},\lambda)\}$
2:procedure ClassifyAndExtract($I,E,LLM$)
3:     $\Xi\leftarrow\emptyset$
4:     for all $i\in I$ do
5:         $p\leftarrow$ CreatePrompt($i,E$)
6:         $response\leftarrow$ QueryLLM($p,LLM$)
7:         $(t_{y},\lambda)\leftarrow$ ParseResponse($response$)
8:         $\Xi\leftarrow\Xi\cup\{(i,t_{y},\lambda)\}$
9:     end for
10:     return $\Xi$
11:end procedure
12:procedure CreatePrompt($i,E$)
13:     $p\leftarrow``\quad"$
14:     for all $e\in E$ do
15:         $p\leftarrow p+\text{``{Example:}"}$: ”
16:         $p\leftarrow p+\text{``{Intent:} "}+e[0]$
17:         $p\leftarrow p+\text{``{Type:} "}+e[1]$
18:         $p\leftarrow p+\text{``{Keywords:} "}+e[2]$
19:     end for
20:     $p\leftarrow p+\text{``{New Intent:} "}+i$
21:     $p\leftarrow p+\text{``{Type, Keywords}"}$ ”
22:     return $p$
23:end procedure
24:procedure QueryLLM($p,LLM$)
25:     return LLM.Query($p$)
26:end procedure
27:procedure ParseResponse($response$)
28:     $type\leftarrow$ ExtractType($response$)
29:     $keywords\leftarrow$ ExtractKeywords($response$)
30:     return $(type,keywords)$
31:end procedure
This intent processing functionality is presented as a User Interface (UI) and intent translation application in Fig. 1. The job of this application is to work as a UI for a human operator and process the intents with crucial information to feed the intent validation and the meta-controller application. We further elaborate our intent processing methodology via Algorithm 1. The approach can be categorized as prompt engineering, where the prompt serves as a template for the LLM to infer the correct response based on minimal data examples.
IV-B Predictive Intent Validation Using Transformer